# The CASE Framework - A New Architecture for Participatory Research and Digital Health Surveillance

### v2

@@ Abstract.
We present the CASE framework, an open-source platform for adaptive, context-aware participatory research, and pandemic preparedness. CASE implements an event-driven architecture that enables dynamic survey workflows, allowing real-time adaptation based on participant responses, external data, temporal conditions, and evolving user states. The framework supports a broad range of research needs, from simple one-time questionnaires to complex longitudinal studies with advanced conditional logic.
Built on over a decade of practical experience, CASE underwent a major architectural rework in 2024, transitioning from a microservice-based design to a streamlined monolithic architecture. This evolution significantly improved maintainability, flexibility, and accessibility to deployment, particularly for institutions with limited technical capacity.
@@
This paper describes the foundations of CASE, details its architectural evolution, and presents lessons learned from real-world deployments. We establish CASE as a mature and reusable research infrastructure that balances sophisticated functionality with practical implementation, addressing the critical global need for sustainable and institutionally controlled data collection systems.
@@ 1. Introduction
The CASE framework introduces a new approach within digital survey methodology, designed to overcome these limitations through an innovative approach to survey design and execution. Unlike conventional survey tools that rely on static branching logic, CASE implements an event-driven architecture that enables dynamic and context-aware workflows. Rather than treating surveys as fixed sequences of questions, CASE models them as responsive systems that can adjust in real time based on multiple inputs, such as participant responses, temporal conditions (even ensuring minimal time distances to validate triggers), external data, or changes in user state. The state of a participant is persistently tracked by the platform and can be updated by the same triggers that influence the progression of the survey. This allows flows to diverge or iterate based on real-time conditions and makes CASE well suited for longitudinal studies and research scenarios where flexibility and contextual awareness are important.
@@
This paper makes three key contributions. First, we present the design and implementation of an event-driven architecture tailored for adaptive participatory research, enabling context-sensitive real-time survey workflows. Second, we document the architectural evolution of the CASE framework, from an initial microservice-based design to a simplified, maintainable monolithic architecture based on insights gained through real-world deployments. Third, we share lessons learned from five years of real-world deployments across diverse research domains and regulatory environments, including national surveillance platforms, specialized health monitoring systems, and cross-domain applications, providing practical insight for future participatory research infrastructure. These deployments, which involve tens of thousands of participants, also serve to validate the scalability and adaptability of the framework.
The remainder of this paper is organized as follows. Chapter 2 outlines the foundation and history of the framework and discusses related work. Chapter 3 provides an overview of the key modules and features that make up the CASE framework. These are the result of extensive collaboration with domain experts over several years, with input from a variety of stakeholders influencing the evolution of the framework. Chapter 4 details the architectural improvements that started in 2024, explaining the motivations behind the rework and technical choices. and adaptions are laid out. Chapter 5 presents selected real-world deployments and lessons learned from these implementations, providing practical insights for future participatory research infrastructure.
@@ 2.1. Background
The first implementation of those concepts began back in 2018 and was driven by a practical need to update the aging Influenzanet platform. This participatory surveillance system for influenza-like illness (ILI) had been in operation since 2003. Over the years, it has expanded into a network of national platforms in 10 European countries, collectively engaging tens of thousands of volunteers each season, providing valuable data for epidemiological research and disease monitoring (Koppeschaar et al., 2017; Paolotti et al., 2014). However, by the mid-2010s, the underlying software of the platform was becoming outdated and increasingly difficult to maintain. It was not designed to incorporate new diseases or modern data practices, such as smartphone input or flexible consent models or enhanced privacy protection. CASE took the role as the successor to this legacy software as a next generation framework to carry forward the citizen science concept of Influenzanet on a more robust, scalable, and sustainable technology.
@@ 2.2. Related Work
Most mainstream survey platforms like LimeSurvey, SurveyMonkey, or Qualtrics provide a web-based questionnaire design with conditional branching and basic logic elements. LimeSurvey offers an expression manager for defining boolean conditions (LimeSurvey GmbH, 2025), while Qualtrics supports complex survey flows with visibility logic and embedded data integration (Qualtrics, 2025). These tools improve on classic paper surveys by allowing skip patterns and piping of answers into subsequent questions. However, they remain fundamentally linear with fixed tree structures, where branches depend only on previous responses. Integration of real-time external context is extremely limited and is typically restricted to preloading known data. Dynamic adaptation during ongoing surveys (e.g., altering question flow based on external input or evolving conditions) is not supported, especially when it involves participant context or state. Although relatively user-friendly and sufficient for basic research, they lack the event-driven architecture and live context integration needed for highly adaptive or longitudinal study designs. More sophisticated platforms like Alchemer support API integration and custom scripting. However, they are still rely on predefined survey structures and do not support a full reconfiguration of the survey logic in response to live context or changes in participant state (Alchemer LLC, 2025). Basic tools like Google Forms and Microsoft Forms offer simple survey creation, but lack advanced logic capabilities, and are primarily designed for casual data collection rather than research applications.
@@
Researchers have explored context-aware survey frameworks that react to sensor data or environmental events. Intille proposed early concepts of context-aware experience sampling (Intille et al., 2003), while the MyExperience framework provided systems to trigger in situ questionnaires based on sensor readings (Froehlich et al., 2007). Studies demonstrated that mobile sensors could initiate relevant questions at opportune moments, for example, asking a user when the accelerometer detects physical activity or when the GPS indicates arrival at a location, improving data relevance and accuracy (Srinivas et al., 2019). In the initial phase of CASE development, the use of smartphone sensors was explored to enhance participatory research capabilities (Hirsch et al., 2018), capturing movement patterns and environmental context to augment traditional self-reported responses. However, such systems, while widely explored (Cornet and Holden, 2018), remained largely proof-of-concept rather than deployed platforms. Implementation required significant custom programming and tight coupling of survey logic with mobile apps. Technical challenges increased as mobile operating systems introduced stricter privacy controls and background processing restrictions (Petter et al., 2019). Earlier context-aware systems focused on momentary interactions rather than providing infrastructure for multi-year or large cohort studies. Recent research continues to seek adaptive survey solutions, including data-driven survey generation approaches (Velykoivanenko et al., 2024), highlighting the ongoing demand for platforms that intelligently adapt to the context of the user. However, a gap remains between research prototypes and robust general-purpose frameworks for longitudinal studies.
@@ 4.2. Disadvantages of the Microservice Approach
These challenges are not limited to CASE (Krug et al., 2024) but indicate the need for a reevaluation of the system architecture, considering the actual requirements and usage patterns of the participatory system.
@@ 5.1. Selected Research Projects
Currently, the ongoing VERDI project (SARS-coV2 variants Evaluation in pRegnancy and paeDIatrics cohorts, since 2021, Horizon Europe, 30+ partners) [^3] VERDI https://cordis.europa.eu/project/id/101045989 uses CASE as a reference use case to explore how participatory surveillance systems can support pandemic preparedness. VERDI is investigating the technological and methodological foundations required for resilient, scalable, and engaging data collection infrastructures that can quickly adapt to newly emerging infectious disease threats.
@@ 5.2. Selected Use Cases
For flu and COVID-19 surveillance, the infectieradar.nl (Netherlands) and flusurvey.net (UK) platforms, both part of the Influenzanet network, demonstrate the capabilities of the CASE framework for large-scale participatory disease monitoring. These platforms engage tens of thousands of volunteers in weekly symptom reporting to track influenza, COVID-19, and other infectious diseases (McDonald et al., 2021; Koppeschaar et al., 2017). In addition to seasonal surveillance, they also support ad hoc one-time surveys hosted within the same application ecosystem. More European platforms within the Influenzanet network have successfully transitioned from legacy technical infrastructure to CASE-based implementations, including grippenet.fr (France), influweb.org (Italy) and grippenet.ch (Switzerland). While the open-source nature of CASE enables such migrations, not all implementations are publicly documented. These platforms operate independently with their own customized implementations of the CASE framework, tailored to their specific national contexts and requirements.
@@
The Post-COVID Research Portal at postcovidonderzoek.nl serves as a national entry point for individuals with long-term symptoms after COVID-19 infection to participate in ongoing scientific studies (National Institute for Public Health and the Environment , Netherland(2025), RIVM). Developed from the start with the reworked version of the CASE software, the platform collects baseline and follow-up data while enabling centralized participant recruitment across multiple coordinated research projects. By streamlining intake and periodic follow-up engagement, the system improves long-term cohort management and enables targeted substudies. The platform also supports integration with clinical studies, as participants are invited through the portal to participate in, for instance, blood bank studies and random control trials, illustrating how CASE allows the digital research infrastructure to connect efficiently with laboratory workflows (UMC Utrecht, 2025).
@@ 6. Closing Remarks and Next Steps
In this paper, we introduce the CASE framework, a modern, flexible, and open-source architecture for the deployment of adaptive surveys, grounded in more than a decade of participatory surveillance experience. Designed to meet the demands of public health infrastructure, CASE has demonstrated robustness in large-scale real-world deployments, including national COVID-19 monitoring platforms and longitudinal post-infection cohort studies. At the same time, CASE is well suited for general-purpose use. Its modular event-driven logic, privacy-aware design, and highly customizable workflows make it suitable for behavioral science, environmental health, and other domains where context-aware participatory data collection is critical.
The 2024 rework ensures that CASE remains aligned with modern standards and research needs, significantly enhancing both maintainability and deployment flexibility. The new architecture retains flexibility and scalability, ensuring that the system remains relevant and extendable for future applications. By combining technical robustness with a user-friendly experience for participants and ethical design, CASE establishes itself as a research infrastructure capable of supporting both urgent epidemiological surveillance and long-term interdisciplinary studies. Indeed, the recent adoption of CASE by numerous European platforms, particularly within the Influenzanet network, highlights its growing role as a foundational infrastructure for multi-study participatory research.
@@
Despite these strengths, several limitations remain. The lack of formal usability evaluations limits the understanding of participant and researcher experience across diverse deployments. Although the monolithic redesign improved maintainability, a technical barrier still exists for non-specialist teams. Furthermore, the long-term sustainability of the framework depends on continued institutional support and the development of a formal governance and funding strategy to ensure ongoing maintenance and community growth.
To address these challenges, we now aim to expand accessibility by enhancing the core tool set, further generalizing CASE’s application beyond public health, making it quicker to deploy, and enhancing preparedness and research reproducibility across various domains.
To maximize the framework’s impact and accessibility, future plans consider a library of ready-to-deploy application templates tailored to common research scenarios, possibly complemented by establishing a Software-as-a-Service (SaaS) platform that will provide hosted CASE instances with managed infrastructure and support services. This would significantly reduce the technical barrier to entry, allowing researchers to rapidly deploy sophisticated participatory systems without requiring extensive technical expertise or infrastructure investment.

### v3

@@ Abstract.
We present CASE, an open-source framework for adaptive participatory research and disease surveillance. Unlike traditional survey platforms with static branching logic, CASE uses an event-driven architecture that adjusts survey workflows in real time based on participant responses, external data, temporal conditions, and evolving participant state. This design supports everything from simple one-time questionnaires to complex longitudinal studies with sophisticated conditional logic.
Built on over a decade of practical experience, CASE underwent major architectural changes in 2024. We replaced a complex microservice design with a streamlined monolithic architecture, significantly improving maintainability and deployment accessibility, particularly for institutions with limited technical resources.
@@
This paper describes the foundations of CASE, documents its architectural evolution, and shares lessons learned from real-world deployments across diverse research domains and regulatory environments. We position CASE as a mature research infrastructure that balances sophisticated functionality with practical deployment needs for sustainable and institutionally controlled data collection systems.
@@ 1. Introduction
CASE addresses these limitations through an event-driven architecture where surveys respond dynamically to multiple inputs rather than following predetermined paths. Unlike conventional survey tools that rely on static branching logic, CASE implements an event-driven architecture that enables dynamic and context-aware workflows. Rather than treating surveys as fixed sequences of questions, CASE models them as responsive systems that can adjust in real time based on multiple inputs, such as participant responses, temporal conditions (even ensuring minimal time distances to validate triggers), external data, or changes in user state. CASE tracks each participant’s state continuously. The same triggers that advance a survey, like a submitted response, a timer event, or external data, can also update participant state. Survey flows can therefore diverge or iterate based on real-time conditions, making CASE suitable for longitudinal studies where context matters.
@@
This paper makes three key contributions. First, we present the design and implementation of an event-driven architecture tailored for adaptive participatory research, enabling context-sensitive real-time survey workflows. Second, we document the architectural evolution of the CASE framework, from an initial microservice-based design to a simplified, maintainable monolithic architecture based on insights gained through real-world deployments. This includes development of graphical configuration tools that enable study creation without programming expertise. Third, we share lessons learned from five years of real-world deployments across diverse research domains and regulatory environments, including national surveillance platforms, specialized health monitoring systems, and cross-domain applications, providing practical insight for future participatory research infrastructure. These deployments, which involve tens of thousands of participants, also serve to validate the scalability and adaptability of the framework.
The remainder of this paper is organized as follows. Section 2 outlines the foundation and history of the framework and discusses related work. Section 3 provides an overview of the key modules and features that make up the CASE framework. These are the result of extensive collaboration with domain experts over several years, with input from a variety of stakeholders influencing the evolution of the framework. Section 4 details the architectural improvements that started in 2024, explaining the motivations behind the rework and the resulting technical choices and adaptations. Section 5 presents selected real-world deployments and lessons learned from these implementations, providing practical insights for future participatory research infrastructure.
@@ 2.1. Background
The first implementation of those concepts began back in 2018 and was driven by a practical need to update the aging Influenzanet platform. This participatory surveillance system for influenza-like illnesses (ILI) has been in operation since 2003. Over the years, it has expanded into a network of national platforms in 10 European countries, collectively engaging tens of thousands of volunteers each season, providing valuable data for epidemiological research and disease monitoring (Koppeschaar et al., 2017; Paolotti et al., 2014). However, by the mid-2010s, the underlying software of the platform was becoming outdated and increasingly difficult to maintain. It was not designed to incorporate new diseases or modern data practices, such as smartphone input or flexible consent models or enhanced privacy protection. CASE took the role as the successor to this legacy software as a next generation framework to carry forward the citizen science concept of Influenzanet on a more robust, scalable, and sustainable technology.
@@ 2.2. Related Work
Most mainstream survey platforms like LimeSurvey, SurveyMonkey, or Qualtrics provide a web-based questionnaire design with conditional branching and basic logic elements. LimeSurvey offers an expression manager for defining boolean conditions (LimeSurvey GmbH, 2025), while Qualtrics supports complex survey flows with visibility logic and embedded data integration (Qualtrics, 2025). These tools improve classic paper surveys by allowing skip patterns and piping of answers into subsequent questions. However, they remain fundamentally linear with fixed tree structures, where branches depend only on previous responses. Integration of real-time external context is extremely limited and is typically restricted to preloading known data. Dynamic adaptation during ongoing surveys (e.g., altering question flow based on external input or evolving conditions) is not supported, especially when it involves participant context or state. Although relatively user-friendly and sufficient for basic research, they lack the event-driven architecture and live context integration needed for highly adaptive or longitudinal study designs. More sophisticated platforms like Alchemer support API integration and custom scripting. However, they still rely on predefined survey structures and do not support a full reconfiguration of the survey logic in response to live context or changes in participant state (Alchemer LLC, 2025). Basic tools like Google Forms and Microsoft Forms offer simple survey creation, but lack advanced logic capabilities, and are primarily designed for casual data collection rather than research applications. These limitations become more apparent in mobile and field contexts, where researchers need offline capability and regulatory compliance.
@@
Researchers have explored context-aware survey frameworks that react to sensor data or environmental events. Intille proposed early concepts of context-aware experience sampling (Intille et al., 2003), while the MyExperience framework provided systems to trigger in situ questionnaires based on sensor readings (Froehlich et al., 2007). Studies demonstrated that mobile sensors could initiate relevant questions at fitting moments, for example, asking a user when the accelerometer detects physical activity or when the GPS indicates arrival at a location, improving data relevance and accuracy (Srinivas et al., 2019). In the initial phase of CASE development, the use of smartphone sensors was explored to enhance participatory research capabilities (Hirsch et al., 2018), capturing movement patterns and environmental context to augment traditional self-reported responses. However, such systems, while widely explored (Cornet and Holden, 2018), remained largely proof-of-concept rather than deployed platforms. Implementation required significant custom programming and tight coupling of survey logic with mobile apps. Technical challenges increased as mobile operating systems introduced stricter privacy controls and background processing restrictions (Petter et al., 2019). Earlier context-aware systems focused on momentary interactions rather than providing infrastructure for multi-year or large cohort studies. Recent research continues to seek adaptive survey solutions, including data-driven survey generation approaches (Velykoivanenko et al., 2024), highlighting the ongoing demand for platforms that intelligently adapt to the context of the user. However, a gap remains between research prototypes and robust general-purpose frameworks for longitudinal studies.
@@ 3.3.4. Messaging System
3.3.5. Configuration Tools
Early versions of CASE required researchers to manually code survey definitions and study logic, creating barriers for teams without programming expertise. To address this, we developed graphical configuration tools that enable study creation and management through visual interfaces. These tools were developed as part of ongoing platform improvements and are publicly available as part of the CASE framework.
Survey Editor
Survey creation previously required manual coding of survey definitions but can now be managed through a graphical interface with structured item editing and a built-in simulator. The Survey Editor allows researchers to construct surveys by arranging items hierarchically, like questions, display text, instructional content, and groups without writing code.
Key features include:
  • Visual item arrangement with drag-and-drop reordering
  • Multilingual content management for international studies
  • Built-in simulator for testing survey behavior before deployment
  • Real-time preview showing how surveys will appear to participants
Expression Editor
Rules and display conditions can now be created with a visual tool rather than manual scripting, speeding up iteration. The Expression Editor represents logical expressions as nested functions in a graphical interface. Researchers configure conditions, validation rules, and dynamic behavior by selecting functions and setting parameters visually.
For example, to show a follow-up question only when a participant selected ”Other” in a previous dropdown, the researcher selects ”Response contains any of these keys” and visually specifies which question and option to check. The tool generates the underlying expression automatically. This approach makes complex conditional logic accessible to researchers without programming background.
Admin UI
A unified web interface consolidates study configuration, participant management, messaging templates, and access control. The Admin UI provides direct access to the Survey Editor, Expression Editor, and other configuration tools through a card-based dashboard (Figure 4). It integrates all study management functions into a single interface, reducing the need to interact with multiple systems or command-line tools.
Figure 4. CASE Admin UI main menu, providing access to study configuration, participant management, messaging and editors for surveys and expressions. This interface addresses the need for non-technical teams to configure studies without manual coding.
Screenshot of the CASE Admin UI showing a modern web interface with a card-based layout. The main screen displays six tools: Study configurator, Participants, User management, Messaging, Standalone editors, and Documentation. Each is represented by an icon and brief description.
While these tools significantly lower technical barriers for study configuration and management, deployment and initial platform setup still require institutional IT support. The tools focus on the operational phase, like creating studies, managing participants, configuring surveys rather than infrastructure deployment.
@@ 4.2. Disadvantages of the Microservice Approach
These challenges emerged during sustained operation across multiple production deployments with different institutional partners, rather than from purely theoretical considerations. They are also not limited to CASE (Krug et al., 2024) but indicate the need for a reevaluation of system architectures based on the actual requirements and usage patterns of participatory research infrastructures.
@@ 5.1. Selected Research Projects
The VERDI project (SARS-coV2 variants Evaluation in pRegnancy and paeDIatrics cohorts, since 2021-2025, Horizon Europe, 30+ partners) [^3] VERDI https://cordis.europa.eu/project/id/101045989 uses CASE as a reference infrastructure to investigate how participatory surveillance systems can support pandemic preparedness under real-world constraints. VERDI is investigating the technological and methodological foundations required for resilient, scalable, and engaging data collection infrastructures that can quickly adapt to newly emerging infectious disease threats.
@@ 5.2. Selected Use Cases
For flu and COVID-19 surveillance, the infectieradar.nl (Netherlands) and flusurvey.net (UK) platforms, both part of the Influenzanet network, demonstrate the capabilities of the CASE framework for large-scale participatory disease monitoring. These platforms engage tens of thousands of volunteers in weekly symptom reporting to track influenza, COVID-19, and other infectious diseases (McDonald et al., 2021; Koppeschaar et al., 2017). In addition to seasonal surveillance, they also support ad hoc one-time surveys hosted within the same application ecosystem.
Table 1 presents selected CASE deployments with verified participant statistics, demonstrating the framework’s capacity to support sustained citizen engagement across different research contexts. Weekly participant numbers reflect 2024-2025 influenza season averages from Influenzanet.info data explorer (ISI Foundation, 2025). Tekenradar reports weekly participants and yearly incident counts (van Vliet and Bron, 2025; National Institute for Public Health and the Environment , Netherlands(2025b), RIVM). The RTR participant count from 2021 and 2025 German federal election study combined (Maier et al., 2022; RPTU University Kaiserslautern-Landau - Political Communication Research Group, 2025). Additional CASE-based platforms operate in Belgium, Switzerland, Italy, Spain, Estonia, and Portugal but current participation statistics were not publicly available at time of writing.
Table 1. Selected CASE-based participatory surveillance deployments showing continuous operational scale.
Platform Country Active Participants
Infectieradar NL 13,700/week (peak: 37,700 in 2020)
Grippenet FR 3,400/week
Flusurvey UK 1,900/week
Tekenradar NL 350/week (6,000 tick bites in 2024)
RTR App DE 2,500 (debates 2021+2025 combined)
Beyond these documented deployments, additional European platforms within the Influenzanet network have successfully transitioned from legacy technical infrastructure to CASE-based implementations, including infectieradar.be (Belgium), gripiradar.ee (Estonia), grippenet.fr (France), influweb.org (Italy), gripenet.pt (Portugal), epidemiradar.es (Spain), and grippenet.ch (Switzerland). While the open-source nature of CASE enables such migrations, not all implementations provide participation statistics or are publicly documented. These platforms operate independently with customized implementations of the CASE framework, tailored to their specific national contexts and requirements.
@@
The Post-COVID Research Portal at postcovidonderzoek.nl serves as a national entry point for individuals with long-term symptoms after COVID-19 infection to participate in ongoing scientific studies (National Institute for Public Health and the Environment , Netherlands(2025a), RIVM). Developed from the start with the reworked version of the CASE software, the platform collects baseline and follow-up data while enabling centralized participant recruitment across multiple coordinated research projects. By streamlining intake and periodic follow-up engagement, the system improves long-term cohort management and enables targeted substudies. The platform also supports integration with clinical studies, as participants are invited through the portal to participate in, for instance, blood bank studies and random control trials, illustrating how CASE allows the digital research infrastructure to connect efficiently with laboratory workflows (UMC Utrecht, 2025).
@@
5.3. Validation Through Deployment
We validate CASE through sustained real-world operation rather than controlled experiments. This approach fits infrastructure research where robustness, maintainability, and institutional fit matter more than isolated performance metrics.
The deployments in Section 5.2 provide evidence along several dimensions:
  • Scale and duration: Multiple platforms have operated continuously for several years, collectively engaging tens of thousands of participants, most of them on regular basis (weekly or quarterly).
  • Adaptability: The same code-base supports influenza surveillance, post-COVID cohort studies, and real-time political sentiment analysis—demonstrating domain-agnostic design.
  • Institutional adoption: Multiple European countries migrated from legacy Influenzanet infrastructure to CASE, suggesting the framework meets institutional requirements for data sovereignty, regulatory compliance, and long-term maintainability.
  • Development sustainability: The 2024 architectural rework was motivated by multi-year operation challenges. That we could perform this rework while maintaining backward compatibility demonstrates architectural resilience.
Part of validation occurred within EU-funded research projects (Section 5.1) focused on pandemic preparedness. Work within these multinational consortia revealed practical constraints around rapid redeployment, cross-country operation, and long-term sustainability under changing conditions. These constraints motivated the architectural changes described in Section 4 and development of graphical configuration tools (Section 3.3.5) that enable research teams to create and adapt studies without programming expertise.
In this context, preparedness means more than epidemiological readiness, it refers to the ability of research infrastructures to be rapidly configured, deployed, and adapted, particularly during emerging health threats when external dependencies may be unreliable.
Beyond high-level dimensions, we can trace specific technical capabilities to deployment contexts. Table 2 provides this mapping.
Table 2. Mapping of CASE capabilities to evidence from real-world deployment contexts (Section 5.2).
Capability Evidence
Adaptive study workflows Used in longitudinal surveillance and cohort follow-up settings .
Context-aware survey logic Individualized follow-ups and state-dependent interactions in production systems.
Longitudinal participant state management Required for sustained cohort participation and multi-study engagement.
Integration of temporal and rule-based logic Used for scheduled interactions and state-dependent follow-ups in deployed systems.
Operational maintainability Architectural redesign motivated by multi-year operation across multiple platforms.
This validation approach does not provide controlled comparisons with alternative platforms, formal usability metrics, or quantified measures of configuration effort. Future work should include systematic usability studies and reproducible benchmarks for common participatory surveillance workflows.
@@ 5.4. Lessons Learned
Reducing Technical Barriers Through Tooling
The developed graphical configuration tools address a critical preparedness challenge: enabling rapid study deployment by teams without programming expertise. Survey creation previously required manual coding of survey definitions, but researchers can now build surveys through the visual Survey Editor with structured item editing and real-time preview. Similarly, study logic and display conditions can be created with the Expression Editor instead of manual scripting, speeding up iteration during time-sensitive deployments. While these tools significantly improve accessibility, infrastructure deployment still requires IT support, highlighting the ongoing tension between usability and technical complexity in participatory research platforms.
@@ 6. Closing Remarks and Next Steps
We presented CASE, an open-source framework for participatory adaptive research built on over a decade of disease surveillance experience. Designed to meet the demands of public health infrastructure, CASE has demonstrated robustness in large-scale real-world deployments, including national COVID-19 monitoring platforms and longitudinal post-infection cohort studies. Additionally, CASE is well suited for general-purpose use. Its modular event-driven logic, privacy-aware design, and highly customizable workflows make it suitable for behavioral science, environmental health, and other domains where context-aware participatory data collection is critical.
The 2024 rework ensures that CASE remains aligned with modern standards and research needs, significantly enhancing both maintainability and deployment flexibility. The new architecture retains flexibility and scalability, ensuring that the system remains relevant and extendable for future applications. By combining technical robustness with a user-friendly experience for participants and ethical design, CASE establishes itself as a research infrastructure capable of supporting both urgent epidemiological surveillance and long-term interdisciplinary studies. The recent adoption of CASE by numerous European platforms, particularly within the Influenzanet network, highlights its growing role as a foundational infrastructure for multi-study participatory research.
@@
Several limitations remain. We lack formal usability evaluations of participant and researcher experience across deployments. The monolithic redesign improved maintainability, and we developed graphical configuration tools (Survey Editor, Expression Editor, Admin UI) that enable researchers to create studies without coding. However, initial deployment and infrastructure setup still require IT expertise, limiting accessibility for some institutions. Most critically, long-term sustainability depends on continued institutional support and formal governance, not just technical design. Without sustainable funding for documentation, maintenance, and institutional handover, even successful platforms risk obsolescence between crises.
To address these challenges, we will consolidate and stabilize existing tooling rather than adding new features. In practice, this means improving documentation, simplifying configuration workflows, and further refining the user-facing tools that already support study setup and deployment. Our aim is not to remove flexibility, but to make common workflows easier to understand, reproduce, and adapt by teams without extensive technical expertise.
Based on repeating patterns observed across deployments, we also plan to provide a small set of example applications that can serve as starting points for common participatory research scenarios, such as longitudinal cohort studies or weekly surveillance reports. Rather than acting as fixed templates, these examples are intended to capture design decisions and best practices that have proven useful in practice. In parallel, we are exploring whether a hosted deployment model could be useful for some organizations without requiring technical expertise or proper in-house infrastructure, while still preserving CASE’s emphasis on self-hostability and data sovereignty.